OpenStudy (anonymous):

how do I do these summations again? Find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . .

OpenStudy (perl):

what is the difference between two consecutive terms?

OpenStudy (anonymous):

n-5?

OpenStudy (perl):

should be just -5

OpenStudy (perl):

so 1 + -5 = -4 -4 + -5 = -9 -9 + -5 = -14 , etc

OpenStudy (anonymous):

yes I got that part:) how do I set up the summation calculation for this though?

OpenStudy (perl):

there is a formula for the sum of an arithmetic sequence

OpenStudy (perl):

first you have to find the 12th term

OpenStudy (amistre64):

we dont really need the 12th term do we? i mean its fine ... but theres a rewriting of it in terms of a1 that suffices

OpenStudy (amistre64):

an = a1 + d(n-1) a1 + a12 = a1 + a1+d(12-1)

OpenStudy (perl):

to find a12 you can use the sequence formula an = a1 + (n-1)*d

OpenStudy (amistre64):

just a little trivia is all \[\frac{n}{2}(2a_1+d(n-1))=\frac{n}{2}(a_1+a_n)\]

OpenStudy (perl):

thats a simpler formula :)

OpenStudy (amistre64):

algebrating it ... \[n~a_1 + \frac{d}{2}(n-1)\] tomato tomato lol

OpenStudy (anonymous):

sorry I was in the bathroom:p but I got n=2 a1=1 and d=-5 but what do I do with the summation sigma thing? I forgot how to set it up. I added the first 12 numbers physically and I got -318

OpenStudy (perl):

so we derived a formula for you Sn = n/2 ( 2*a1 + d(n-1) )

OpenStudy (perl):

S12 = 12/2 ( 2 * 1 + (-5)*(12-1) )

OpenStudy (anonymous):

ahh ok that works to. is there a way to solve it using the sigma sign tho? \[\sum_{n=?}^{n}\]

OpenStudy (amistre64):

if we know the rule that defines the sequence .. \[\sum_{n=1}^{k} a_n\] \[\sum_{n=1}^{k} d(n-1)+a_1\] \[\sum_{n=1}^{k} dn-d+a_1\] \[d\sum_{n=1}^{k} n+(a_1-d)\sum_{n=1}^{k}\] and this requires still knowing how to sum up a sequence if consecutive integers to produce \[d\frac{1+k}{2}+k(a_1-d)\]

OpenStudy (anonymous):

so what exactly would I plug into the calculator?